You want to purchase 50 shares of IBM stock at 80 from your broker. The initial margin is 50%

a. Construct your balance sheet when you open the position at P=$80.

b. suppose the stock price falls to 50, construct the balance sheet at that price

c. calculate the percentage margin at P=$50. If the maintenance margin is 30%, will you receive a margin call>

d. suppose you close your position at p=50. what is your return from this transaction?
Answer #1

Solution:)

A) No of shares = 50

Price = 80

Initial Margin =50%

Total Investment= 50 * 80 * 50% = 4000 *.5 = 2000

Balance Sheet

Both the entries take place on the asset side i.e.

Cash goes down by 2000

And Investment/Tradeable securities increases by 2000

2) If the stock price dec to 50, still the balance sheet remains the same as the unrealized gain from the trading securities goes into the Income statement but no effect on the balance sheet.

3) Percentage Margin at P=50

=> (80-50)/80 = 30/80 = .375 = 37.5% loss in value

We know the margin call is triggered at price => P* (1-Initial Margin)/(1 - Maintenance Margin)

=> 50*(1 - 50%)/ (1 - 30%) = 50*(1-.5)/(1-.3) = 50*(.5/.7) = 50*.714 = 35.714

4) Now if the position is closed at p=50

Loss = Net loss * Leverage ratio(Since we used leverage in the form of IM)

= ( (80 -50) / 80 ) * (1/IM)

= .375 * (1/0.5) = .375 *2 = .75 = 75% loss on the given transaction.